why are electrons negatively charged?

Grade:11

3 Answers

RAGHVAN DAS
12 years ago

Well, you could give a limited explanation by regarding protons and electrons from the quark level.
On this level, the explanation of the difference in the electron/proton charge (and, for that matter, the neutron's non-charge) is that electrons and protons are constituted from different sets of quarks (i.e. quarks with different charges)
I believe that the proton is constituted by 2 quarks with 2/3 charge units each, along with one quark that has -1/3 charge unit (2/3+2/3-1/3=1).
The electron, I believe, consists of 3 -1/3-quarks (-1/3-1/3-1/3=-1)
The neutron (2/3-1/3-1/3=0) (I think..)

Only thing I can think of to explain it is: that's the way the universe is. The particles have a property that attracts and repels based on something which we have arbitrarily defined as a "charge". The "charge" behaves according to certain rules, and it is best described in terms of positive and negative terms.

Jaya IITK
12 years ago

 dont thnk theres an answer at this level..

 

u need 2 get higher knwledge

 

BEST OF LUCK!!!!!

kanika mathur
12 years ago

william crookes studied the electrical conductivity of gases . where he found that a ray  was emitted out from cathode known as cathode rays..

when these rays were subjected to electric field , cathode rays was deflected towards positive plate and .so they were characterised as rays which contain certain negatively charged particles called electrons .

i hope u like the answer ....